OpenStudy (anonymous):

Find the fifth term in this sequence: t sub n = 12 minus 2 n

OpenStudy (anonymous):

If I'm getting this right, the sequence is : Tn= 12 - 2*n. The first term is T1 and to find it, you replace the n=1 in the right side of the equation - which means : T1= 12 - 2=10 Same goes for T2, which is : T2=12-2*2=8 Ergo, the fifth term is T5, which is : T5=12-2*5=2.